The lenght of the garden is 6 meters and 58 meters of encing is required what is the width of the garden
Answer by umail08(87)

Let: L=Length and W=Width
Perimeter = 2L + 2W
58 = 2(6) + 2W
58 = 12 + 2W
58 = 12 + 2W
58 -12 = 2W -12 + 12
46 = 2W
46/2 = 2W/2
23 = W